OpenStudy (anonymous):

can someone help me please ? 3. Given 4x – 8y = 8: a. Transform the equation into slope-intercept form. b. Find the slope and y-intercept of the line. c. Find the equation, in point-slope form, of the line that is perpendicular to this line and passes through the point (1, 2).

OpenStudy (nurali):

or the line with the equation below: a.Transform the equation into slope-intercept form. 4x–8y=8 8y=4x-8 y=x/2-1 .. b.Find the slope and y-intercept of the line. Standard form of equation of a line:y=mx+b, m=slope, b=y-intercept For given equation:4x–8y=8 slope=1/2 y-intercept=-1 .. c.Find the equation, in point-slope form, of the line that is perpendicular to this line and passes through the point (1, 2). slope of perpendicular line=-2 (negative reciprocal of slope of given line) point-slope form of equation: y1-y2=m(x1-x2) using coordinates of given point (1,2), Equation: y-2=-2(x-1)

OpenStudy (mrnood):

@nurali do you think that giving the answer in so much detail is likely to encourage the poster to LEARN how to do this. It is better to guide, and allow the poster to DO some of the steps in order to understand better

OpenStudy (anonymous):

-8y = -4x +8 y = 1/2x +1 b/ y= mx+b slope = 1/2 intercept = 1 ......................................... 3c perpendicular slope is the reciprocal y= mx +b 2 = -2/1(1) +b 2 +2 = b 4= b y= -2x + 4